when the rock is thrown until it hits the street? 2.37 A juggler throws a bowling pin straight up with an initial speed of 8.20 m/s. How much time elapses until the bowling pin returns to the jugglers hand? つ20
Answer #1

Given,

u (initial velocity) = 8.20 m/s

v (final velocity) = 0 m/s (on reaching max. height)

using first equation of motion v = u + at

here a = - 9.8 m/s2

0 = 8.20 - (9.8*t)

t = (8.20/9.8) s

t = 0.84 second

Total time elapsed when it reaches to juggler's is 2 times t (for reaching the max. height for upward direction

and reaching downward direction)

Total time = 2*t

= 2*0.84

= 1.68 second